### v1

# Energy Optimized Piecewise Polynomial Approximation Utilizing Modern Machine Learning Optimizers*

Abstract
This work explores an extension of Machine Learning (ML)-optimized Piecewise Polynomial (PP) approximation by incorporating energy optimization as an additional objective. Traditional closed-form solutions enable continuity and approximation targets but lack flexibility in accommodating complex optimization goals. By leveraging modern gradient descent optimizers within TensorFlow, we introduce a framework that minimizes total curvature in cam profiles, leading to smoother motion and reduced energy consumption for input data that is unfavorable for sole approximation and continuity optimization. Experimental results confirm the effectiveness of this approach, demonstrating its potential to improve efficiency in scenarios where input data is noisy or suboptimal for conventional methods.
Index Terms:
Piecewise Polynomials, Gradient Descent Optimizers, Approximation, TensorFlow, Electronic Cams, Energy Consumption
I INTRODUCTION
I-A Motivation and Problem Statement
PPs are commonly utilized in various scientific and engineering disciplines. One key area of interest is trajectory planning for machines in mechatronics, computing time-dependent position, velocity, or acceleration profiles that provide setpoints for controllers of industrial machines. In this context, electronic cams define the repetitive motion of servo drives, commonly defined as input point clouds and approximated by PPs. Traditionally, the approximation process involves solving a system of equations in closed form, incorporating domain-specific constraints like continuity, cyclicity, or periodicity. While computationally efficient, closed-form solutions offer limited flexibility regarding polynomial degrees, bases, or the integration of additional constraints.
As optimization goals grow more complex, numerical methods such as gradient descent become beneficial. Since gradient-based optimization lies at the heart of training neural networks, modern ML frameworks like TensorFlow or PyTorch come with a range of state-of-the art optimizers. In previous work [5] [1], we utilize these optimizers directly to fit $C^{k}$-continuous PPs for the use in the electronic cam domain. The problem formulation allows for the addition of further domain-specific optimization targets, however, it is not clear whether results converge without introducing further regularization measures.
This is why, in this work, we investigate optimizing the total curvature of PP cam curves as an additional domain-specific goal. This holds the potential of leading to smoother motion profiles and therefore improved efficiency in machines in cases where the input data is noisy or unfavorable for continuity and approximation optimization alone, leading to reduced forces and energy consumption in such scenarios.
I-B Contribution
Considering the $3$-objective optimization target described in section II-B, the set of Pareto efficient solutions for input data unfavorable for sole continuity and approximation optimization can be sketched as depicted in fig. 1. This concerns input data where results end up in the left upper corner of feasible solutions. In this work, we demonstrate that an extension of the framework introduced in [5] and [1] towards energy optimization is feasible and indeed improves results in the case of such input data.
Figure 1: Sketch of Pareto frontier (dashed) for this $3$-objective optimization. Feasible solutions form the cross-hatched area.
I-C Prior and Related Work
Related work in the cam approximation domain mostly either relies on closed-form solutions, like [2], or utilizes parametric functions for approximation of given input data, like B-Spline or NURBS curves, as in [3]. The first is limited with respect to complexity and number of optimization goals and the latter cannot directly be utilized in industrial servo drives utilizing non-parametric PP functions. Our current work represents a natural progression from the principles introduced in [5] and [1].
II PROBLEM FORMULATION
II-A Model Definition
Input data consists of $n$ samples $x_{1}\leq\dots\leq x_{n}\in\mathbb{R}$ with $y_{i}\in\mathbb{R}$. We define the polynomial boundaries of a PP function $f$ on the interval $I=[x_{1},x_{n}]$ as $\xi_{0}\leq\dots\leq\xi_{m}$, with $\xi_{0}=x_{1}$ and $\xi_{m}=x_{n}$. Each of the $m$ segments $p_{i}\colon I\to\mathbb{R}$ is modeled via polynomials of degree $d$ as
$p_{i}=\sum_{j=0}^{d}\alpha_{i,j}x^{j}.$ (1)
The coefficients $\alpha_{i,j}$ are the model parameters trained via TensorFlow gradient descent optimizers according to the loss function described in the next section.
II-B Loss Function
In order to reduce the forces applied to the cam-follower system, we want to optimize the cam’s acceleration curve. To do so, we can quantify total curvature as $\int_{I}f^{\prime\prime}(x)^{2}\;\mathrm{d}{x}\,$ for a cam profile $f$ over the interval $I$. Considering our individual polynomial segments $p_{i}$, we receive
$\int_{I}f^{\prime\prime}(x)^{2}\;\mathrm{d}{x}\,$ $=\sum_{i=1}^{m}\int_{I_{i}}p^{\prime\prime}_{i}(x)^{2}\;\mathrm{d}{x}\,$ (2)
with
$\int_{I_{i}}p^{\prime\prime}_{i}(x)^{2}\;\mathrm{d}{x}\,=\int_{I_{i}}\left(\sum_{j=2}^{d}\alpha_{i,j}\cdot j(j-1)x^{j-2}\right)^{2}\;\mathrm{d}{x}\,.$ (3)
Plugging in the boundary points for the definite integral and generalizing to the overall PP with $m$ segments then leaves us with the energy loss
$\ell_{E}=\sum_{i=1}^{m}\sum_{j=2}^{d}\sum_{k=2}^{d}\alpha_{i,j}\alpha_{i,k}\frac{jk(j-1)(k-1)}{j+k-3}\cdot\left(\xi_{i}^{j+k-3}-\xi_{i-1}^{j+k-3}\right).$ (5)
We quantify the approximation error $\ell_{2}$ as
$\ell_{2}=\frac{1}{n}\sum_{i}|f(x_{i})-y_{i}|^{2}$ (6)
and sum up discontinuities at all $\xi_{i}$ across relevant derivatives for $C^{k}$-continuity as
$\ell_{\text{CK}}=\frac{1}{m-1}\sum_{i=1}^{m-1}\sum_{j=0}^{k}\left(\frac{\Delta_{i,j}}{r_{k}}\right)^{2}\text{with}\Delta_{i,j}=p^{(j)}_{i+1}(\xi_{i})-p^{(j)}_{i}(\xi_{i})$ (7)
and achieve cyclicity by adapting $m-1$ to $m$ and generalizing $\delta_{i,j}=p^{(j)}_{1+(i\bmod m)}(\xi_{i\bmod m})-p^{(j)}_{i}(\xi_{i})$, ignoring the case $j=0$ when $i=m$. We then can prioritize either of the three objectives by defining the weighted total loss as
$\ell=\alpha\ell_{\text{CK}}+\beta\ell_{2}+(1-\alpha-\beta)\ell_{\text{E}},$ (8)
where $\alpha,\beta\geq 0$ and $\alpha+\beta\leq 1$.
III EXPERIMENTAL RESULTS
In an effort of shifting non-$\ell_{\text{E}}$-optimized results to the upper left corner in the sense of fig. 1, we define $100$ data points $x\in[0,1]$ and let $y=\sin{(x^{2}4\pi)}$. We then add noise to the data via NumPy’s random.normal function with a scale of $0.1$. In practice, such data could stem from measurements, introducing some noise to the input cam data. We utilize TensorFlow optimizers to fit our introduced PP model to this data. We run $1000$ epochs with early stopping and patience of $100$ epochs.
To have enough free parameters, we fit a $C^{2}$-continuous PP of degree $7$. In addition to the nature of the data, in an effort of further shifting results to the upper left corner in the sense of fig. 1, each segment should not have too many data points, i.e. the number of polynomial segments needs to be sufficiently high. Table I shows the result of fitting a $16$ segment PP with and without $\ell_{\mathrm{E}}$ optimization utilizing the AMSGrad version of Adam in this context.
TABLE I: Results for fitting a $C^{2}$-continuous PP of degree $7$.
Parameters $\ell_{2}$ $\ell_{\mathrm{CK}}$ $\ell_{\mathrm{E}}$
$\alpha=0.9,\beta=0.1$ $0.0079$ $0.0026$ $5.3551$
$\alpha=0.9,\beta=0.099$ $0.0096$ $0.0210$ $0.9326$
Table I confirms that optimizing $\ell_{\text{E}}$ in addition to $\ell_{2}$ and $\ell_{\text{CK}}$ works. Oscillations in the curve’s shape are reduced, therefore providing a smoother result. This can be observed from plots of the resulting PPs along with further experiments available at [4].
IV SUMMARY AND OUTLOOK
Our results demonstrate that an extension of the framework introduced in [5] and [1] towards energy optimization is feasible. Results converge without additional regularization measures and indeed improve in the case of input data that is unfavorable for sole approximation and continuity optimization.
Further work could focus on hyperparameter optimization or measures to improve performance of $\ell_{\text{E}}$ optimization in combination with orthogonal bases.